John is twice as old as his friend Peter. Peter is 5 years older than Alice. In 5 years, John will be three times as old as Alice. How old is Peter now?
Found 2 solutions by dpaulin, dmf30633:
Answer by dpaulin(12)

lots of possible ways to find this answer I honestly am not sure which is best I just use trail and error. first I think of all the possible statements
Peter x 2 = john
John/2 = peter
peter -5 = alice
alice + 5 = peter
alice x 3 = john
john/3 = alice
now that I have that in my head, I ask my self what age would john be when alice was just born. if you force yourself to do a problem like this in your head it will help you in the future. think about the possibilities will if john is 10 divide that by 2 which mean peters is 5 subtract 5 and alice was just born she is 0. now in 5 years alice will be 5 and john 15 divided john 3 it that proves alice is 5. add 5 to alice and we get john who is now 10 which is the answer.
**note it may be easier for you to start with the youngest of the age group and work from there.

Answer by dmf30633(3)

John is twice as old as his friend peter, J=2P
Peter is 5 years older than Alice, P=A+5
In 5 years John will be three times as old as Alice, J+5=3A

J=2P
P=A+5
J+5=3A
-------------
J=2(A+5)
J=2A+10
-------------
J+5=3A
2A+10+5=3A
10+5=A
15=A
-------------
P=A+5
P=15+5
P=20
So peter is 20, lets check our work. If peter is 20 then John must be 40. In 5 years John will be 45, which is 3 times as old as Alice who is 15.